OpenStudy (anonymous):

what is y=1/8x+8 in standard form

OpenStudy (texaschic101):

standard form is : Ax + By = C......A needs to be positive, and there can be no fractions. y = 1/8x + 8 -- subtract 1/8x from both sides -1/8x + y = 8 -- got to get rid of the fraction, so we will multiply by the common denominator which is 8 8(-1/8x + y = 8) = -x + 8y = 64 -- now we have to make x positive, so we multiply by -1 -1(-x + 8y = 64) = x - 8y = -64 <=== standard form